OpenStudy (anonymous):

The displacement (in meters) of a particle moving in a straight line is given by s=t^2-5t+16. Find the average velocity of the time interval (3,4) please help. no idea

OpenStudy (ajprincess):

Let the time interval be (x,y). Then the average velocity during that time interval will be v=(s_y-s_x)/(y-x) v=(s_4-s_3)/(4-3) s_4=4^2-5(4)+16 s_3=3^2-5(3)+16

OpenStudy (ajprincess):

Am i right @hartnn?

hartnn (hartnn):

i didn't know about this method, i would have used calculus...(integration)...don't know whether the asker can/should use calculus.

OpenStudy (ajprincess):

hmm k. i too got to know this method recently only.

OpenStudy (anonymous):

Yes it is for calc

OpenStudy (anonymous):

I need to solve it in terms of calculus but I don't know how to get the answer

hartnn (hartnn):

do u know the relation between displacement and velocity?

OpenStudy (anonymous):

well just the formula f(a+h)-f(a)/h that's all I have

OpenStudy (anonymous):

no not at all, but this a question for my homework that is due tonight

OpenStudy (anonymous):

R u there?

hartnn (hartnn):

yup. that formula is similar to one used by ajprincess in first comment, could you understand that?

OpenStudy (anonymous):

not exactly. I understand in terms of (x,y) but have no idea what she is doing with it. Perhaps you can explain what to do please if its not to much?

hartnn (hartnn):

sure. u have V= s(t+h)-s(t)/h where V is the average velocity and s(t) =s=t^2-5t+16. now you have interval (3,4) , so we put t=3, h=1 we get V= (s(4)-s(3))/(1) Did u get this much ??

hartnn (hartnn):

ask if there is any doubt in any step

OpenStudy (anonymous):

yes I got that far I am just checking my algebra. Why did you put h=1?

hartnn (hartnn):

h= time interval = 4-3 = 1

hartnn (hartnn):

or you can say change in time = 4-3=1

OpenStudy (anonymous):

ok thank you. I got it. So for the answer did you get 2 or am I totally wrong?

hartnn (hartnn):

what did u get s(4) and s(3) as ?

hartnn (hartnn):

oh, yes even i am getting + its correct :)

hartnn (hartnn):

+2

OpenStudy (anonymous):

yes 2 right? So basically I am plugging those numbers instead of t right?

hartnn (hartnn):

yes, s(4) means put t= 4 in s(t)

hartnn (hartnn):

and remember you are also dividing by time interval

OpenStudy (anonymous):

yes but h will still be 1 right?

OpenStudy (anonymous):

since 4-3=1

hartnn (hartnn):

in this problem, yes, h= 1

hartnn (hartnn):

any more queries ?

OpenStudy (anonymous):

I got it Thank you so much for your help. I really appreciate it

hartnn (hartnn):

welcome ^_^ and \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\)

OpenStudy (anonymous):

Thank you!

OpenStudy (ajprincess):

Sorry cud nt explain u my answer. had to go in the middle:(. Thanx @hartnn for explaining it. great work @hartnn:)